OpenStudy (freemap):

A line is parallel to y = 1/3 x + 5 and passes through (–6, 0). A. y = –3x B. y = –3x + 2 C. y = 1/3 x + 2 D. y = 1/3 x – 6 Which is the equation of the line?

OpenStudy (madhu.mukherjee.946):

option C

OpenStudy (madhu.mukherjee.946):

do i need to explain

OpenStudy (freemap):

Thanks, yes please

OpenStudy (madhu.mukherjee.946):

okay

OpenStudy (anonymous):

parallel means same slope

OpenStudy (madhu.mukherjee.946):

see two parallel lines have their slopes equal

OpenStudy (dinamix):

its y=1/3(-6) +2 =0

OpenStudy (freemap):

I know, I thought you had to plug in again

OpenStudy (madhu.mukherjee.946):

the line you have has a slope of 1/3

OpenStudy (madhu.mukherjee.946):

so the required line will also have the same slope

OpenStudy (madhu.mukherjee.946):

do you get?

OpenStudy (freemap):

okay thanks so much

OpenStudy (freemap):

I do get it, computer lagging

OpenStudy (madhu.mukherjee.946):

now u use the formula y-y1=m(x-x1)

OpenStudy (madhu.mukherjee.946):

whenever you are asked to find a slope of a line wich passes through a point u will use this formula

OpenStudy (madhu.mukherjee.946):

now just plug in the values